# Efficient Exact Algorithms for Minimum Covering of Orthogonal Polygons with Squares

### v1

@@ Abstract
The Orthogonal Polygon Covering with Squares (OPCS) problem takes as input an orthogonal polygon without holes where vertices have integral coordinates i.e. a polygon $P$ without any holes, and with $n$ vertices and axis-parallel edges. Moreover, each vertex has integer-valued $x$- and $y$-coordinates. The aim is to find a minimum number of axis-parallel, possibly overlapping squares which lie completely inside $P$, such that their union covers the entire region inside $P$. Aupperle et. al [3] provide an $\mathcal{O}(N^{1.5})$-time algorithm to solve OPCS for orthogonal polygons without holes, where $N$ is the number of integral lattice points lying in the interior or on the boundary of $P$. Moreover, designing algorithms for OPCS with a running time polynomial in $n$ (the number of vertices of $P$) was discussed as an open question in [3], since $N$ can be exponentially larger than $n$. In this paper we design a polynomial-time exact algorithm for OPCS with a running time of $\mathcal{O}(n^{14})$.
We also consider the following structural parameterized version of the problem. A knob in an orthogonal polygon is a polygon edge whose both endpoints are convex polygon vertices. Given an input orthogonal polygon without holes that has $n$ vertices and $k$ knobs, we design an algorithm for OPCS with running time $\mathcal{O}(n^{2}+k^{14}\cdot n)$. This algorithm is more efficient than the above exact algorithm when $k=o(n^{13/14})$.
In [3], the Orthogonal Polygon with Holes Covering with Squares (OPCSH) problem is also studied. Here, the input is an an orthogonal polygon which could have holes and the objective is to find a minimum square covering of the input polygon. This version where the orthogonal polygon is allowed to have holes is shown to be NP-complete (even when all lattice points in the interior or on the boundary of the polygon constitute the input). We think there is an error in the existing proof in [3], where a reduction from Planar 3-CNF is shown. We fix this error in the proof with an alternate construction of one of the gadgets used in the reduction, hence completing the proof of NP-completeness of OPCSH, when all lattice points in the interior or on the boundary of the polygon are given as input.
@@ 1 Introduction
We consider the problem of covering orthogonal polygons with a minimum number of (possibly overlapping) axis-parallel squares.
An orthogonal polygon is a simple polygon such that every polygon-edge is parallel to either the x-axis or the y-axis (Figure 1(a)). We wish to cover a given orthogonal polygon with a minimum number of (possibly overlapping) axis-parallel squares, i.e. given an orthogonal polygon $P$ we wish to find a minimum number of squares such that every square lies inside the region of $P$ and the union of the regions of the plane covered by these squares is the region of the plane covered by polygon $P$ itself (Figure 1(b) and Figure 1(c)).
@@
In general, an orthogonal polygon $P$ may contain holes. However, for most of the paper, we deal with problems where the input is an orthogonal polygon without any holes. First, we define the problem of Orthogonal Polygon Covering with Squares (OPCS) formally as follows.
@@
Note that for this problem, it is possible that the input polygon is given in terms of its $n$ vertices and their coordinates. Since the polygon is such that all vertices have integral coordinates, it is also possible that the input polygon is given in terms of its boundary vertices as well as all lattice points contained inside $P$. Let this set of lattice points lying on the boundary of $P$ or inside $P$ be of size $N$. Note that there could be an exponential factor difference between the values of $N$ and $n$.
We also study a parameterized version of OPCS, called $p$-OPCS. Before we define the problem, let us consider a knob as a structure occurring in an orthogonal polygon $P$. A knob is an edge where both endpoints are convex vertices of $P$ (the formal definition of knobs is given in Section 2).
@@
The parameterized problem of $p$-OPCS takes the number of knobs in the input orthogonal polygon to be a structural parameter, to be utilized when designing parameterized algorithms.
@@
Once again, the polygon $P$ could be described in terms of its $n$ boundary vertices, or in terms of the $N$ lattice vertices that lie on the boundary of $P$ or inside $P$.
@@
In [3], exact polynomial time solutions of OPCS are discussed, where the input is considered to be the set of all $N$ lattice points inside (in the interior and the boundary of) $P$. The algorithms crucially use following associated graph $G(P)$.
@@
Clearly from the definition, we know that any set of blocks covered simultaneously by a single square shall form a clique. Further, [1] shows the converse as well.
@@
In [3], $G(P)$ is shown to be a chordal graph (please refer to Definition 2.13 for the definition of a chordal graph) if $P$ does not have holes. In [3] it is further shown that using the polynomial time algorithm for minimum clique covering in chordal graphs (discussed in [7]), OPCS can be solved in polynomial time. If the number of lattice points inside $P$ is $N$, then the algorithm takes $\mathcal{O}(N^{2.5})$ time. In [3], there is also references provided that would result in an algorithm with complexity $\mathcal{O}(N^{1.5})$ to solve minimum clique cover in the associated chordal graph $G(P)$ using more structural properties of the orthogonal polygon $P$.
@@
Aupperle et. al [3] also prove that OPCSH is NP-Hard if the input orthogonal polygon $P$ with possible holes is described in terms if the $N$ lattice points lying on the boundary of $P$ as well as inside $P$. The paper reduces Planar 3-CNF [8] to this problem in order to prove its NP-Hardness.
Bar-Yehuda [4] describes an $\mathcal{O}(n\log n+\textsc{OPT})$ $2$-factor approximation algorithm to solve this problem where $n$ is the number of vertices and OPT is the minimum number of squares to cover $P$.
@@
Culberson and Reckhow [6] proves that the problem of minimum covering of orthogonal polygons with rectangles is NP-hard. Kumar and Ramesh [2] provide a polynomial time approximation algorithm with approximation factor of $\mathcal{O}(\sqrt{\log N})$ for the problem of covering orthogonal polygons (with or without holes) with rectangles. Most of the other geometric minimum cover problems for orthogonal polygons are NP-complete as well [10].
Our Results.
@@
In the results due to [3], the input size is considered to be the number of lattice points inside $P$ or lying on the boundary of $P$. However, it is inefficient to represent orthogonal polygons with all the lattice points inside it. Rather, it would be more efficient to take an orthogonal polygon $P$ as a sequence of the boundary vertex points of $P$ (in either clockwise or counter-clockwise direction). In case $P$ has one or more holes, then this kind of input lists the sequence of vertices on the outer boundary, followed by the sequence of vertices on the boundary of each hole of $P$.
@@
As an example, if $P$ is itself a large square with vertices $(0,0)$, $(0,2^{t})$, $(2^{t},0)$, $(2^{t},2^{t})$, then even though it can be represented by just $O(t)$ bits by simply listing down the vertices in order, we would need $\Omega(4^{t})$ bits to actually represent all lattice points that lie inside or on the boundary of $P$. Moreover, if OPT is the actual minimum number of squares to completely cover $P$, $OPT$ can also potentially be exponential in the number of bits needed to represent the vertices (e.g. $P$ is a $1\times 2^{t}$ rectangle).
@@
In this paper we consider the input size $n$ as the number of polygon vertices of $P$ and then design efficient algorithms for minimum square covering with respect to $n$. Note that the algorithm in [3] becomes exponential now, as there can be exponentially many points inside $P$. Designing algorithms which are polynomial in the number $n$ of polygon vertices was stated as an open question in [3]. To the best of our knowledge, this is the first study to answer these questions. This algorithm with a running time of $\mathcal{O}(n^{14})$, where $n$ is the number of vertices of the input polygon, and related structural results are given in Section 3 and Section 4. Moreover, our algorithm also works for polygons with rational coordinates as we can simply scale up the polygon by an appropriate denominator without changing the number of vertices $n$.
@@
In Section 5, we consider the $p$-OPCS problem and further optimize the above algorithm, considering the polygon to have $n$ vertices and $k$ knobs (edges with both endpoints subtending $90^{\circ}$ to the interior of the polygon, formally defined in Definition 2.1). We design a recursive algorithm running in time $\mathcal{O}(n^{2}+k^{14}\cdot n)$. This algorithm is more efficient than the above exact algorithm if $k=o(n^{13/14})$.
Finally in Section 6, we fix an error in the proof discussed in [3] of NP-completeness of minimum square covering of orthogonal polygons with holes. Note that such a result implies that we do not expect an algorithm for OPCSH that has a running time polynomial in the number $n$ of boundary vertices of the input orthogonal polygon $P$ which may have holes.
@@ 2 Preliminaries
We denote by $P$ our input polygon with $n$ vertices. In the entire paper, we always consider the vertices of the polygon to have integral coordinates. Unless mentioned otherwise, the polygon $P$ is assumed to not have any holes. Let the vertices of $P$ in order be $v_{1},v_{2},\ldots,v_{n}$, where $v_{i}=(x_{i},y_{i})$ with $x_{i}$ and $y_{i}$ being integer coordinates. For convenience, we define $v_{0}:=v_{n}$ and $v_{n+1}:=v_{1}$. Therefore $(v_{i},v_{i+1})$ is a polygon edge of $P$ for all $i\in\{1,2,\ldots,n\}$. For any natural number $d$, we denote the set $\{1,2,\ldots,d\}$ by the notation $[d]$. Next, let the minimum number of squares required to cover $P$ be denoted by $\textsc{{OPCS}}(P)$. We use the terms left, right, top, bottom, vertical, horizontal to indicate greater x-coordinate, smaller x-coordinate, greater y-coordinate, smaller y-coordinate, parallel to y-axis, and parallel to x-axis, respectively. By boundary of a square (rectangle), we mean its four sides. A block lying on the boundary of a square is one which lies inside the square and where the boundary of the block overlaps with the boundary of the square. We assume that arithmetic and logic operations on two numbers take constant time. For any two points $a$, $b$ on the plane, we denote the (Euclidean) length of the line segment joining $a$ and $b$ as $\overline{ab}$.
{observation}
Given an input polygon $P$ of OPCS, the minimum number of squares to cover $P$ remains the same, even if $P$ is scaled up by an integral factor and/or rotated $90^{\circ}$ (clockwise/counter-clockwise).
Let us formally define convex and concave vertices of an orthogonal polygon.
Definition 2.1 (Concave and Convex vertices of $P$).
@@
For the OPCS problem, we are only interested in maximal squares that are completely contained in the region of the input polygon.
@@
An axis parallel square $S$ is called a valid covering square of of an orthogonal polygon $P$ or simply a valid square if $S$ is fully contained in the region of $P$.
@@
A valid square $S$ is a maximal valid square or simply a maximal square, if no other valid square with a larger area contains $S$ entirely(Figure 3).
@@
Note that for a given minimum square covering of an input orthogonal polygon $P$, some squares may be maximal squares (Definition 2.2) while the other squares may be valid squares which are not maximal. However, for each such valid square $S$ which is not maximal, we can replace it with any maximal square containing $S$ (this exists as $S$ is not maximal and the interior of $P$ is a bounded region with finite area) to obtain another minimum square covering of $P$. Since this process does not change the count of the squares used to cover $P$ and since the newly introduced squares cover the entire area of the removed squares, the new covering is a minimum covering with maximal squares. {observation} There exists a minimum square covering of any orthogonal polygon $P$ where every square of the covering is a maximal square.
The paper [3] defines a structure called knobs. We define knobs slightly differently but they essentially are the same.
@@
A knob in an orthogonal polygon $P$ is an edge $(v_{i},v_{i+1})$ of $P$ such that $v_{i}$ and $v_{i+1}$ are both convex vertices of $P$ (Figure 4(a)). Suppose $v_{i}$ and $v_{i+1}$ share the same x coordinate. Then we say that they form a left knob if $(v_{i},v_{i+1})$ is a left boundary edge (i.e. all points just to the left of the edge are in the exterior of $P$), otherwise they form a right knob. Similarly, when $v_{i}$ and $v_{i+1}$ share the same y coordinate, then we say that they form a top knob if $(v_{i},v_{i+1})$ is a top boundary edge (i.e. all points just to the top of the edge are in the exterior of $P$), otherwise they form a bottom knob.
@@
Given an orthogonal polygon $P$, a convex vertex $v_{i}$ is said to be a non-knob convex vertex if neither $(v_{i},v_{i+1})$ nor $(v_{i-1},v_{i})$ is a knob (i.e. both $v_{i-1}$ and $v_{i+1}$ are concave vertices).
We further introduce strips as follows.
@@
A strip of an orthogonal polygon $P$ is a maximal axis-parallel non-square rectangular region $Y$ lying inside $P$ such that each side in the longer set of parallel sides of $Y$ is contained entirely in a polygon edge of $P$ (Figure 4(b)).
@@
Given an orthogonal polygon $P$, for any pair of parallel polygon edges $e_{1}$, $e_{2}$, there can be at most one strip with its longer sides lying inside $e_{1}$ and $e_{2}$ respectively.
@@
Given an orthogonal polygon $P$, a rec-pack of $P$ is an axis parallel rectangle $R$ (of dimension $t\times\eta t$ or $\eta t\times t$) with integer aspect ratio $\eta$ (i.e. ratio of the lengths of the longer side to its shorter side) that lies completely inside $P$ (Figure 5(a)). We define the length $t$ of the shorter side of the rec-pack $R$ to be the width of $R$ and the aspect ratio $\eta$ to be the strength of $R$.
@@
A valid square is a rec-pack of strength $1$. We consider any square to be a trivial rec-pack.
For a rec-pack $R$ with width $t$ and strength $\eta$, it can be covered with exactly $\eta$ many $t\times t$ valid squares. We formally define this operation in the next definition.
@@
Given an orthogonal polygon $P$ and a rec-pack $R$ of width $t$ and strength $\eta$ in $P$, we define the extraction of $R$, $\mathcal{E}(R)$ to be the sequence of $\eta$ valid squares of size $t\times t$ placed side by side to cover the entire region of $R$ (Figure 5(b)). Moreover, for a set of rec-packs $\mathcal{R}$, we define its extraction $\mathcal{E}(\mathcal{R})$ to be the union of the extraction of each rec-pack in $\mathcal{R}$.
$\mathcal{E}(\mathcal{R})=\bigcup\limits_{R\in\mathcal{R}}\mathcal{E}(R)$
@@
(b) Extraction $\mathcal{E}(R)$ of rec-pack $R$
@@
For a square $S$ (a trivial rec-pack), the extraction is a singleton set containing $S$ itself, i.e. $\mathcal{E}(S)=\{S\}$. Moreover, for any set $\mathcal{S}$ of squares, we have $\mathcal{E}(\mathcal{S})=\mathcal{S}$.
Note that any side-by-side placed sequence of $\eta$ many $t\times t$ valid squares covers up a rec-pack of width $t$ and strength $\eta$. Therefore, whenever there is a solution with a side-by-side placed sequence of congruent squares, we can keep track of the rec-pack which gives the set of these congruent squares upon extraction. But first, we need to define a covering using rec-packs.
@@
A set of rec-packs $\mathcal{R}$ is said to cover the input orthogonal polygon $P$ if the set of squares in $\mathcal{E}(\mathcal{R})$ cover the polygon $P$. Moreover, we say $\mathcal{R}$ to be a minimum covering, if $\mathcal{E}(\mathcal{R})$ is a minimum cardinality set of squares covering $P$ and no two rec-packs $R_{1},R_{2}\in\mathcal{R}$ have $\mathcal{E}(R_{1})\cap\mathcal{E}(R_{2})\neq\emptyset$ (i.e. produce some common square).
@@
For any set $\mathcal{S}$ of squares covering $P$, $\mathcal{S}$ is also a set of rec-packs covering $P$. Similarly, for any minimum cardinality set of squares $\mathcal{S}$ covering $P$, $\mathcal{S}$ is also a set of rec-packs which is a minimum covering of $P$. Finally, if a set of rec-packs $\mathcal{R}$ is a covering (minimum covering) of $P$, then the set of squares $\mathcal{E}(\mathcal{R})$ is a set of trivial rec-packs which is also a covering (minimum covering).
We also state the following observation.
{observation}
A set of rec-packs $\mathcal{R}$ covers an orthogonal polygon $P$ if and only if all the rec-packs in $P$ together form a set of rectangles covering $P$.
@@
Given a graph $G$, we denote the vertex set by $V(G)$ and the edge set by $E(G)$. The neighbourhood of a vertex $v\in V(G)$ is denoted by $N(v)$. The closed neighbourhood of $v$ is denoted by $N[v]$. Given a graph $G$, a subgraph $H$ is an induced subgraph of $G$ if $V(H)\subseteq V(G)$ and $E(H)$ contains all edges of $E(G)$ such that both endpoints are in $V(H)$. In this case, we also say that the vertex set $V(H)$ induces the subgraph $H$. Recall the definition of chordal graphs and simplicial nodes [5].
@@
We finally reiterate one of the major results of [3].

### v2

# Efficient Algorithms for Minimum Covering of Orthogonal Polygons with Squares

@@ Abstract
Let $P$ be an orthogonal polygon (polygon having axis-parallel edges) of $n$ vertices, without holes. The Orthogonal Polygon Covering with Squares (OPCS) problem takes as input such an orthogonal polygon $P$ with integral vertex coordinates, and asks to find the minimum number of axis-parallel squares whose union is $P$ itself. Aupperle et. al [4] provide an $\mathcal{O}(N^{1.5})$-time algorithm to solve OPCS for orthogonal polygons without holes, where $N$ is the number of integral lattice points lying in the interior or on the boundary of $P$. In their paper, designing algorithms for OPCS with a running time polynomial in $n$, the number of vertices of $P$, was stated as an open question; $N$ can be arbitrarily larger than $n$. Output sensitive algorithms were known due to Bar-Yehuda and Ben-Chanoch [6], but these fail to address the open question, as the output can be arbitrarily larger than $n$. We address this open question by designing a polynomial-time exact algorithm for OPCS with a worst-case running time of $\mathcal{O}(n^{10})$.
We also consider the following structural parameterized version of the problem. Let a knob be a polygon edge whose both endpoints are convex polygon vertices. Given an input orthogonal polygon without holes that has $n$ vertices and at most $k$ knobs, we design an algorithm for OPCS with a worst-case running time $\mathcal{O}(n^{2}+k^{10}\cdot n)$. This algorithm is more efficient than the former, whenever $k=o(n^{9/10})$.
The problem of Orthogonal Polygon with Holes Covering with Squares (OPCSH) is also studied by Aupperle et. al [4]. Here, the input is an orthogonal polygon which could have holes and the objective is to find a minimum square covering it. They claim a proof that OPCSH is NP-complete even when all lattice points inside the polygon constitute the input. We think there is an error in their proof, where an incorrect reduction from Planar 3-CNF is shown. We provide a correct reduction with a novel construction of one of the gadgets, and show how this leads to a correct proof of NP-completeness of OPCSH.
@@ 1 Introduction
An orthogonal polygon is a simple polygon such that every polygon-edge is parallel to either the x-axis or the y-axis (Figure 1(a)). We consider the problem of covering a given orthogonal polygon with the minimum number of (possibly overlapping) axis-parallel squares.
Formally, given an orthogonal polygon $P$, the problem is to find the minimum number of squares such that every square lies inside the region defined by $P$, and the union of the squares is the entire polygon $P$ itself (Figure 1(b) and Figure 1(c)). Equivalently, the problem is to find the minimum number of axis-parallel squares whose union is exactly $P$.
@@
For most of the paper, we deal with orthogonal polygons without holes. We formally define the problem of Orthogonal Polygon Covering with Squares (OPCS).
@@
We also study a parameterized version of OPCS, called $p$-OPCS. Let a knob be an edge where both of its endpoints are convex vertices of $P$ (the formal definition of knobs is given in Section 2). The parameterized problem of $p$-OPCS takes the number of knobs in the input orthogonal polygon to be a structural parameter, to be utilized when designing parameterized algorithms. We define $p$-OPCS as follows.
@@
Note that for these problems, there are a couple of ways to represent the polygon $P$. One way is to just provide the coordinates of the $n$ vertices, in clockwise/counter-clockwise order. Another way is to list down all lattice points lying inside $P$. If there are $N$ lattice points lying inside $P$, $N$ can be even exponential in $n$.
@@
In the works of Aupperle, Conn, Keil and O’Rourke [4], exact polynomial time solutions of OPCS are discussed, where the input is considered to be the set of all $N$ lattice points inside $P$. The algorithms crucially use the notion of associated graph $G(P)$.
@@
By definition, any set of blocks covered by a single square shall correspond to a clique in the associated graph $G(P)$. Further, Albertson and O’Keefe [2] show the converse as well.
@@
In the works of Aupperle et. al [4], $G(P)$ is shown to be a chordal graph, if $P$ does not have holes (please refer to Definition 2.14 for the definition of a chordal graph). They further show that, using the polynomial time algorithm for minimum clique covering in chordal graphs (due to Gavril [12]), OPCS can be solved in polynomial time with respect to the number of lattice points, $N$, inside $P$. An algorithm with running time $\mathcal{O}(N^{2.5})$ is presented and techniques are mentioned to speed it up to $\mathcal{O}(N^{1.5})$, using specific structural properties of the associated graph $G(P)$.
@@
Aupperle et. al [4] also claim a proof that OPCSH is NP-hard if the input orthogonal polygon $P$ (possibly with holes) is described in terms of the $N$ lattice points inside $P$. The NP-complete problem Planar 3-CNF [15] is reduced to OPCSH in order to prove its NP-hardness.
The works of Bar-Yehuda and Ben-Chanoch [6] consider the problem of OPCS where the input is the $n$ vertices of the polygon $P$. They provide an exact algorithm solving OPCS in $\mathcal{O}(n+\textsc{OPT})$ time and $\mathcal{O}(n)$ space, where OPT is the minimum number of squares to cover $P$. This is particularly interesting as this eliminates the dependence on $N$, which could be arbitrarily larger than $n$. They crucially use the concept of essential squares and build up the solution one square at a time, to finally get an optimal set of squares. Since they considered arbitrary placement of vertices (not necessarily at lattice points), the concept of associated graphs was irrelevant to their approach. However, in the worst case OPT can be arbitrarily larger than $n$; for example, a $1\times t$ rectangle has $n=4$ and $\text{OPT}=t$. Hence this algorithm, or any other output-sensitive algorithm, can never provide a worst-case running time guarantee of a polynomial in $n$. Later, in another paper, Bar-Yehuda [5] describes an $\mathcal{O}(n\log n+\textsc{OPT})$ $2$-factor approximation algorithm to solve this problem where $n$ is the number of vertices and OPT is the minimum number of squares to cover $P$. Incidentally the same paper states that no polynomial time exact algorithm with respect to $n$ and OPT were known.
@@
The works of Culberson and Reckhow [10] prove that the problem of minimum covering of orthogonal polygons with rectangles is NP-hard. Kumar and Ramesh [3] provide a polynomial time approximation algorithm with an approximation factor of $\mathcal{O}(\sqrt{\log N})$ for the problem of covering orthogonal polygons (with or without holes) with rectangles. Most of the other geometric minimum cover problems for orthogonal polygons are NP-complete as well [17]. Another closely related problem is the tiling problem, where the squares additionally need to be non-overlapping [1]. This is already non-trivial when the input is a rectangle [13, 19].
Our results and the organization of the paper.
@@
In the algorithms due to Aupperle et. al [4], the input size is considered to be the number of lattice points $N$ lying inside $P$. However, this is an inefficient way to represent orthogonal polygons. Rather, it would be more efficient to represent an orthogonal polygon $P$ as a sequence of its vertices (in either clockwise or counter-clockwise direction). This is because a polygon with $n$ vertices can have arbitrarily large number of lattice points $N$ inside it [6].
@@
Even if we consider the total number of bits required to represent the vertices to be $n^{\prime}$, then $N$ can be exponential in $n^{\prime}$. As an example, consider $P$ to be a large square with vertices $(0,0)$, $(0,2^{t})$, $(2^{t},0)$, $(2^{t},2^{t})$. Even though it can be represented by just $\mathcal{O}(t)$ bits, there are $\Theta(4^{t})$ lattice points inside it. Moreover, if OPT is the minimum number of squares required to completely cover $P$, then recall that OPT can also potentially be exponential in the number of bits needed to represent the vertices (for example $P$ being a $1\times 2^{t}$ rectangle).
@@
In this paper we consider the input to be the $n$ vertices of the polygon $P$, and we design efficient algorithms for OPCS, with respect to $n$. The algorithm by Aupperle et. al [4]. becomes exponential now, as there can be exponentially many lattice points inside $P$. In their paper, designing algorithms which are polynomial in the number $n$ of polygon vertices was stated as an open question. Note that output sensitive algorithms [6] are also exponential in the worst case.
To the best of our knowledge, our paper is the first to answer this open question. We present an algorithm with a running time of $\mathcal{O}(n^{10})$, where $n$ is the number of vertices of the input polygon; along with related structural results, in Section 4 and Section 5. It is interesting to note that our algorithm can also work for polygons with rational coordinates, as we can simply scale up the polygon by an appropriate factor without changing the number of vertices $n$.
@@
In Section 6, we consider the $p$-OPCS problem and further optimize the above algorithm. We consider the polygon to have $n$ vertices and at most $k$ knobs (edges with both endpoints subtending $90^{\circ}$ to the interior of the polygon, formally defined in Definition 2.2). We design a recursive algorithm running in time $\mathcal{O}(n^{2}+n\cdot k^{10})$. This algorithm is more efficient than the former whenever $k=o(n^{9/10})$.
Moreover, we think that the claimed proof of NP-hardness of OPCSH by Aupperle et. al [4] is incorrect, as they incorrectly reduce from a polynomial-time solvable variant of a problem. In Section 7, we provide a correct proof of NP-hardness of OPCSH. Our proof relies on some of the existing constructions [4], but we also provide novel constructions for some gadgets used in the reduction. Note that such a result implies that we do not expect an algorithm for OPCSH that has a running time polynomial in $n$ as well as $N$, as the reduction creates an instance with $N=\Theta(n)$.
@@ 2 Preliminaries
We denote by $P$ our input polygon with $n$ vertices. For distinction, we associate the terms ‘vertices’ and ‘edges’ to the polygon, the terms ‘corners’ and ‘sides’ to other geometric objects, and the terms ‘nodes’ and ‘arcs’ to graphs. In the entire paper, we always consider the vertices of the polygon to have integral coordinates. Unless mentioned otherwise, the polygon $P$ is assumed to not have any holes. We denote the number of lattice points inside $P$ by $N$. We assume that all arithmetic operations take constant time.
Let the vertices of $P$ in order be $v_{1},v_{2},\ldots,v_{n}$, where $v_{i}=(x_{i},y_{i})$ with $x_{i}$ and $y_{i}$ being integer coordinates. For convenience, we define $v_{0}:=v_{n}$ and $v_{n+1}:=v_{1}$. Therefore $(v_{i},v_{i+1})$ is a polygon edge of $P$ for all $i\in\{0,1,2,\ldots,n\}$. For any natural number $d$, we denote the set $\{1,2,\ldots,d\}$ by the notation $[d]$. Let the minimum number of squares required to cover $P$ be denoted by $\textsc{{OPCS}}(P)$. For any two points $a$, $b$ on the plane, we denote the (Euclidean) length of the line segment joining $a$ and $b$ as $\overline{ab}$.
We use the terms left, right, top, bottom, vertical, horizontal to indicate greater x-coordinate, smaller x-coordinate, greater y-coordinate, smaller y-coordinate, parallel to y-axis, and parallel to x-axis, respectively. By boundary of a square (rectangle), we mean its four sides. A block lying on the boundary of a square (rectangle) is one which lies inside the square (rectangle) and where the boundary of the block overlaps with the boundary of the square.
Proposition 2.1.
The minimum number of squares to cover an orthogonal polygon $P$, remains the same when $P$ is scaled up by an integral factor and/or rotated $90^{\circ}$ (clockwise/counter-clockwise).
We formally define convex and concave vertices of an orthogonal polygon.
Definition 2.2 (Concave and convex vertices).
@@
It is interesting to look at maximal squares that are completely contained in the region of the input polygon $P$.
@@
An axis parallel square $S$ is said to be valid if $S$ is fully contained in the region of $P$.
@@
A valid square $S$ is a maximal square, if no other valid square with a larger area contains $S$ entirely (Figure 3).
@@
Note that for a minimum square covering of an orthogonal polygon $P$, some squares may be maximal squares while the other squares may be valid squares which are not maximal. However, for each of such valid square $S$ which is not maximal, we can replace it with any maximal square containing $S$ to obtain another minimum square covering of $P$ containing only maximal squares.
{observation}
There exists a minimum square covering of any orthogonal polygon $P$ where every square of the covering is a maximal square.
We restate the definition of knobs as in previous works [6, 4].
@@
A knob is an edge $(v_{i},v_{i+1})$ of the orthogonal polygon $P$ such that $v_{i}$ and $v_{i+1}$ are both convex vertices of $P$ (Figure 4(a)).
Suppose $v_{i}$ and $v_{i+1}$ share the same x coordinate. Then the edge $(v_{i},v_{i+1})$ is a left knob if it is a left boundary edge (i.e. all points just to the left of the edge are outside $P$), otherwise they form a right knob.
Similarly, when $v_{i}$ and $v_{i+1}$ share the same y coordinate, then the edge $(v_{i},v_{i+1})$ is a top knob if it is a top boundary edge (i.e. all points just to the top of the edge are outside $P$), otherwise they form a bottom knob.
@@
A convex vertex $v_{i}$ is said to be a non-knob convex vertex if neither $(v_{i},v_{i+1})$ nor $(v_{i-1},v_{i})$ is a knob. Equivalently, both $v_{i-1}$ and $v_{i+1}$ are concave vertices.
We introduce a special structure called strips as follows.
@@
A strip of an orthogonal polygon $P$ is a maximal axis-parallel non-square rectangular region $Y$ lying inside $P$, such that each of the longer parallel side of $Y$ is completely contained in a polygon edge of $P$ (Figure 4(b)).
@@
For any pair of parallel polygon edges $e_{1}$, $e_{2}$ of $P$, there can be at most one strip with its sides overlapping with $e_{1}$ and $e_{2}$.
@@
A rec-pack of $P$ is an axis parallel rectangle $R$ of dimension $t\times\eta t$ or $\eta t\times t$ that lies completely inside $P$ (Figure 5(a)). We define the length $t$ of the shorter side of the rec-pack $R$ to be the width of $R$ and the aspect ratio $\eta$ to be the strength of $R$.
@@
A valid square is a rec-pack of strength $1$. Any square is a trivial rec-pack.
For a rec-pack $R$ with width $t$ and strength $\eta$, it can be covered with exactly $\eta$ many $t\times t$ valid squares. We formally define this operation as an extraction.
@@
Given an orthogonal polygon $P$ and a rec-pack $R$ of width $t$ and strength $\eta$, we define the extraction of $R$, $\mathsf{ext}(R)$ to be the set of $\eta$ valid squares of size $t\times t$ placed side by side to cover the entire region of $R$ (Figure 5(b)).
For a set $\mathscr{R}$ of rec-packs, we define its extraction $\mathsf{ext}(\mathscr{R})$ to be the union of the extraction of each rec-pack in $\mathscr{R}$.
$\mathsf{ext}(\mathscr{R})=\bigcup\limits_{R\in\mathscr{R}}\mathsf{ext}(R)$
@@
(b) Extraction $\mathsf{ext}(R)$ of rec-pack $R$
@@
For a square $S$ (a trivial rec-pack), the extraction is a singleton set containing $S$ itself, i.e. $\mathsf{ext}(S)=\{S\}$. Moreover, for any set $\mathscr{S}$ of squares, we have $\mathsf{ext}(\mathscr{S})=\mathscr{S}$.
This structure helps us to efficiently represent a solution of a square covering. If the set of squares constituting the solution has a subset of $\eta$ many side-by-side placed $t\times t$ valid squares, they can be simply replaced by the rec-pack defined by their union (of strength $\eta$ and width $t$).
To exploit this, we need to formally define a covering using rec-packs.
@@
A set of rec-packs $\mathscr{R}$ is said to cover the input orthogonal polygon $P$ if the set of squares $\mathsf{ext}(\mathscr{R})$ cover the polygon $P$. Moreover, $\mathscr{R}$ is said to be a minimum covering, if $\mathsf{ext}(\mathscr{R})$ is a minimum covering of $P$, and no two rec-packs produce the same sqaure on extraction, i.e. for all $R_{1},R_{2}\in\mathscr{R}$, $\mathsf{ext}(R_{1})\cap\mathsf{ext}(R_{2})=\emptyset$.
@@
For any set $\mathscr{S}$ of squares covering $P$, $\mathscr{S}$ is also a set of rec-packs covering $P$. Similarly, for any minimum covering of squares $\mathscr{S}$ of $P$, $\mathscr{S}$ can also be viewed as a set of rec-packs which is a minimum covering of $P$. Similarly, if a set of rec-packs $\mathscr{R}$ is a covering (minimum covering) of $P$, then the set of squares $\mathsf{ext}(\mathscr{R})$ is a set of trivial rec-packs which is also a covering (minimum covering) of $P$.
@@
Given a graph $G$, we denote the node set by $V(G)$ and the arc set by $E(G)$. The neighbourhood of a node $v\in V(G)$ is denoted by $N(v)$. The closed neighbourhood of $v$ is denoted by $N[v]$. Given a graph $G$, a subgraph $H$ is an induced subgraph of $G$ if $V(H)\subseteq V(G)$ and $E(H)$ contains all arcs of $E(G)$ such that both endpoints are in $V(H)$. In this case, we also say that the node set $V(H)$ induces the subgraph $H$. Recall the definition of chordal graphs and simplicial nodes [8].
@@
We finally reiterate one of the major results of the works of Aupperle et. al [4].